Shift summary for eve. 10/7/99

Ed Blucher (blucher@hep.uchicago.edu)
Thu, 7 Oct 1999 23:57:51 -0500


Shift Summary for Eve Shift, 10/7/99

On Shift: Patty McBride, Ed Blucher

Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...):0.5

Runs ended this shift:

14806 Laser scan

========================================================================
Overview of shift:
No beam. TeV purification in progress. Nothing going on at KTeV.

========================================================================
Description of Shift:

21:00 Laser scan

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift